Hi-Tek Manufacturing, Inc., makes two types of industrial component parts—the B300 and the T500. An absorption costing income statement for the most recent period is shown:

Hi-Tek Manufacturing Inc.
Income Statement
Sales $ 1,699,400
Cost of goods sold 1,224,856
Gross margin 474,544
Selling and administrative expenses 630,000
Net operating loss $ (155,456 )

Hi-Tek produced and sold 60,400 units of B300 at a price of $20 per unit and 12,600 units of T500 at a price of $39 per unit. The company’s traditional cost system allocates manufacturing overhead to products using a plantwide overhead rate and direct labor dollars as the allocation base. Additional information relating to the company’s two product lines is shown below:

B300 T500 Total
Direct materials $ 400,700 $ 162,800 $ 563,500
Direct labor $ 120,500 $ 43,000 163,500
Manufacturing overhead 497,856
Cost of goods sold $ 1,224,856

The company has created an activity-based costing system to evaluate the profitability of its products. Hi-Tek’s ABC implementation team concluded that $53,000 and $109,000 of the company’s advertising expenses could be directly traced to B300 and T500, respectively. The remainder of the selling and administrative expenses was organization-sustaining in nature. The ABC team also distributed the company’s manufacturing overhead to four activities as shown below:

Manufacturing
Overhead
Activity
Activity Cost Pool (and Activity Measure) B300 T500 Total
Machining (machine-hours) $ 209,336 90,700 62,100 152,800
Setups (setup hours) 126,420 71 230 301
Product-sustaining (number of products) 101,200 1 1 2
Other (organization-sustaining costs) 60,900 NA NA NA
Total manufacturing overhead cost $ 497,856

Required:

1. Compute the product margins for the B300 and T500 under the company’s traditional costing system.

2. Compute the product margins for B300 and T500 under the activity-based costing system.

3. Prepare a quantitative comparison of the traditional and activity-based cost assignments.
Answer #1

Part A

Predetermined overhead rate = Estimated total manufacturing overhead cost
Estimated total direct labor dollars = 497856/163500 = $3.04 per direct labor dollar

B300

T500

Total

Sales revenue

1208000

491400

1699400

Direct materials

400700

162800

563500

Direct labor

120500

43000

163500

Manufacturing overhead applied (@ $3.04 per direct labor-dollar)

366320

130720

497040

Total manufacturing cost

887520

336520

1224040

Product margin

320480

154880

475360

60400*20 =1208000

12600*39 =491400

Part B

Activity Cost Pools

total cost

total activity

activity rate

machining

209336

152800

1.37

set ups

126420

301

420

product sustaining

101200

2

50600

B300

T500

Total

Sales revenue

1208000

491400

1699400

Direct materials

400700

162800

563500

Direct labor

120500

43000

163500

machining

124259

85077

209336

set ups

29820

96600

126420

product sustaining

50600

50600

101200

Advertising expense

53000

109000

162000

Total manufacturing cost

778879

547077

1325956

Product margin

429121

(55677)

373444

Part C

The quantitative comparison is as follows:

Traditional Cost System

B300

%

T500

%

Total

Direct materials

400700

71.11%

162800

28.89%

563500

Direct labor

120500

73.70%

43000

26.30%

163500

Manufacturing overhead applied (@ $3.24 per direct labor-dollar)

366320

73.70%

130720

26.30%

497040

Total cost assigned to products

887520

336520

1224040

Selling and administrative expense

630000

Total cost

$1854040

Activity-Based Costing System

B300

%

T500

%

Total

Direct costs:

Direct materials

400700

71.11%

162800

28.89%

563500

Direct labor

120500

73.70%

43000

26.30%

163500

Advertising expense

53000

32.72%

109000

67.28%

162000

Indirect costs

machining

124259

59.36%

85077

40.64%

209336

set ups

29820

23.59%

96600

76.41%

126420

product sustaining

50600

50.00%

50600

50.00%

101200

Total cost assigned to products

778879

547077

1325956

Costs not assigned to products:

Selling and administrative

468000

other

60900

Total cost

$1854856

630000-162000 = 468000